# ExploRLLM: Guiding Exploration in Reinforcement Learning with Large Language Models

### 14 Mar 2024 v1

@@ V-3 Code policy generation by LLM
However, as indicated in Figure 2(c), there are instances where the generated affordance map may not be optimal. For example, the optimal pick position for the letter ‘O’ should be at its rim, whereas the heatmap suggests the center. To address those sub-optimum problems, we adopt a stochastic policy based on the affordance map rather than a deterministic policy that selects the point of highest affordance. Given that RL evolves through rewards obtained from environmental interactions, the sub-optimality in exploration policy can be identified and rectified through learning. This approach also offers the potential to provide counter-examples during the phase of replay buffer collection.
@@ VII Results
TABLE I: Comparison of success and error rates of our method against baselines during 50 evaluation episodes with both seen and unseen colors. Evaluation includes different tasks: short-horizon (SH) and long-horizon (LH) and different initialization methods: no overlapping between letters (NO) and allowed overlapping (AO). For the ExploRLLM policy, the standard deviations of 6 seeds are included to show the stability of the training process.
@@
Inner Monologue + CLIPort 0.82 0.72 0.58 0.42 0.20 0.24 0.18 0.26
@@ VII-B RL performance
For short-horizon tasks, as illustrated in Figure I, ExploRLLM shows stable performance, unlike versions without the exploration policy, which sometimes fail to converge and display significant variance in success rates and low-level errors. Our method exceeds LLM-generated policies in success rate, minimizes robot behavior errors, and reduces the gap between NO and AO scenarios, highlighting our exploration policy’s effectiveness in correcting FMs’ inaccuracies. In contrast, the CLIPort-based method struggles with novel scenarios or complex geometric relationships between objects. For long-horizon tasks, RL agents without LLM-based exploitation failed to converge. As depicted in Figure I, ExploRLLM outperforms Socratic Models, Inner Monologue, and the policy generated by LLMs, showcasing superior performance for long-horizon tasks.
VII-C Generalization to unseen long horizon tasks

### 15 Mar 2024 v2

@@ V-3 Code policy generation by LLM
However, as indicated in Figure 2(b), there are instances where the generated affordance map may not be optimal. For example, the optimal pick position for the letter ‘O’ should be at its rim, whereas the heatmap suggests the center. To address those sub-optimum problems, we adopt a stochastic policy based on the affordance map rather than a deterministic policy that selects the point of highest affordance. Given that RL evolves through rewards obtained from environmental interactions, the sub-optimality in exploration policy can be identified and rectified through learning. This approach also offers the potential to provide counter-examples during the phase of replay buffer collection.
@@ VII Results
TABLE I: Comparison of success and error rates of our method against baselines during 50 evaluation episodes with both seen and unseen colors. Evaluation includes different tasks: short-horizon (SH) and long-horizon (LH) and different initialization methods: no overlapping between objects (NO) and allowed overlapping (AO). For the ExploRLLM policy, the standard deviations of 6 seeds are included to show the stability of the training process.
@@
Inner Monologue + CLIPort 0.82 0.72 0.58 0.42 0.18 0.26 0.20 0.24
@@ VII-B RL Performance
For short-horizon tasks, as illustrated in Table I, ExploRLLM shows stable performance, unlike versions without the exploration policy, which sometimes fail to converge and display significant variance in success rates and low-level errors. Our method exceeds LLM-generated policies in success rate, minimizes robot behavior errors, and reduces the gap between NO and AO scenarios, highlighting our exploration policy’s effectiveness in correcting FMs’ inaccuracies. In contrast, the CLIPort-based method struggles with novel scenarios or complex geometric relationships between objects. For long-horizon tasks, RL agents without LLM-based exploitation failed to converge. As depicted in Table I, ExploRLLM outperforms Socratic Models, Inner Monologue, and the policy generated by LLMs, showcasing superior performance for long-horizon tasks.
VII-C Generalization to Unseen Long-horizon Tasks